# Path partial groups

### v5

@@ Abstract.
The first and third authors were partially supported by Ministerio de Ciencia e Innovación (Spain) grant PID2020-118753GB-I00
@@ 1. Introduction
The proof of Theorem A is built upon a functor, defined in Section 3, that maps every pair $(G,\mathcal{H})$, for $G=(V,E)$ a simple undirected graph and =$\mathcal{H}=\{H_{v}\}_{v\in V}$ a collection of non-trivial groups, to a partial group ${\mathbb{M}}(G,\mathcal{H})$ such that the group of automorphisms $\operatorname{Aut}_{{\mathcal{P}}art}\big{(}{\mathbb{M}}(G,\mathcal{H})\big{)}$ is closely related to the groups $\operatorname{Aut}_{{\mathcal{G}}raphs}(G)$ and $\prod_{v\in V}\operatorname{Aut}_{{\mathcal{G}}roups}(H_{v})$ (see Theorem 4.4). The algebraic structure of the partial group ${\mathbb{M}}(G,\mathcal{H})$ is related to path concatenation within $G$. In fact, the special case with $\mathcal{H}$ consisting of copies of ${\mathbb{Z}}_{2}$ is the central character in this work: path partial groups.
@@
The algebraic structure of ${\mathbb{P}}(G)$ can be easily described in terms of paths in $G$, and it resembles the ideas underlying path algebras, i.e., the algebra associated to a quiver [3, II.1.2]. The following result is the key ingredient in the proof of Theorem A.
@@ 2. Basics on partial groups
Let ${\mathcal{M}}$ be a set and $\mathbb{D}\subseteq\mathbb{W}({\mathcal{M}})$ be a subset such that,
@@
    if $u\in\mathbb{D}$ then $(u^{-1},u)\in\mathbb{D}$ and $\Pi(u^{-1}\circ u)=1$.
A word $w\in\mathbb{D}$ will be called non-degenerate if there is no $1$ in $w$. Then, $\mathbb{D}$ is totally determined by its non-degenerated words.
@@
Let $\mathcal{F}(a)=\{1,a,a^{-1}\}$ and set the non-degenerated words of $\mathbb{D}_{a}$ to be all possible words alternating $a$’s and $a^{-1}$’s. In other words, the non-degenerated words of $\mathbb{D}_{a}$ are all the different finite sub-words of the infinite word $(a,a^{-1},a,a^{-1},a,a^{-1},\dots)$. The inversion is understood and, for any word $u\in\mathbb{D}\left(\mathcal{F}(a)\right)$,
@@
An homomorphism $\varphi\colon{\mathcal{M}}_{1}\to{\mathcal{M}}_{2}$ is an isomorphism of partial groups if the map $\varphi$ is bijective and $\varphi^{-1}$ is also a morphism of partial groups. Finally, an automorphism of the partial group ${\mathcal{M}}_{1}$ is an isomorphism $\varphi\colon{\mathcal{M}}_{1}\to{\mathcal{M}}_{1}$.
@@
has alphabet ${\mathcal{M}}$ equal to the quotient set,
${\mathcal{M}}=\cup_{C\in\mathcal{C}}{\mathcal{M}}(C)/\sim$
where $\sim$ is the smallest equivalence relation that contains the following relation,
@@
and such that there is a map $\Pi$ closing the following diagram, where $\mathbb{D}$ is defined as the subset of words $u$ in the free monoid $\mathbb{W}({\mathcal{M}})$ for which there exists an object $C\in\operatorname{Ob}(\mathcal{C})$ and a word $v$ in $\mathbb{D}(C)$ such that $u$ is the image of $v$ under the component-wise application of $\sim$, the rightmost vertical arrow is induced by $\sim$, and the leftmost vertical arrow by its component-wise application,
@@ 3. Partial groups out of simple graphs
In what follows, $G=(V,E)$ is an object in ${{\mathcal{G}}raphs}$ and =$\mathcal{H}=\{H_{v}\}_{v\in V}$ is a collection of groups. Our aim is to define a partial group depending on $G$ and $\mathcal{H}$. Now, for a clique $C\subset V$ in $G$, we consider the free product $H_{C}=\mathop{\ast}\limits_{v\in C}H_{v}$ and the following construction.
@@
    $(u_{1},u_{2},\dots,u_{n})\in\mathbb{D}(\operatorname{CR}(C))\mapsto\Pi(u_{1}\circ u_{2}\circ\cdots\circ u_{n})\in\operatorname{CR}(C)\text{, and}$
@@
$(x_{1},x_{2},\dots,x_{2n}):=(w,w^{-1})=(u_{1},u_{2},\dots,u_{n},u_{n}^{-1},u_{n-1}^{-1},\dots,u_{1}).$
@@
Let $\Delta$ be the poset of cliques $C\subset V$ in $G$. An inclusion $C_{1}\subset C_{2}$ between cliques gives rise to an inclusion map $\operatorname{CR}(C_{1})\subset\operatorname{CR}(C_{2})$ and to its associated map of partial groups ${\mathbb{M}}(C_{1})\to{\mathbb{M}}(C_{2})$. This way, we have a functor ${\mathbb{M}}(-)\colon\Delta\to{{\mathcal{P}}art}$.
@@
Because of the simple shape of the category $\Delta$ and the functor ${\mathbb{M}}(-)$, it is straightforward that, in Definition 3.6, Equation (C1) suffices to define the equivalence relation $\sim$ for the colimit. In fact, we define ${\mathcal{M}}(G,\mathcal{H})$ using only that condition. Then the map
@@
is well defined and injective, where $(g_{1},g_{2},\ldots,g_{r})\in\operatorname{CR}(C)$ for some clique $C$. Moreover, this map together with uniqueness of the reduced form ($\Pi$), ensure that condition (C2) is satisfied too. From now on, we will make the identification given by (S) and the corresponding one for $\mathbb{D}(G,\mathcal{H})$ without further notice.
@@ 4. Graph morphisms from partial groups homomorphisms
The category ${{\mathcal{I}}njPart}$ is the wide subcategory of the category ${{\mathcal{P}}art}$ with morphisms the partial group homomorphisms whose kernel has no torsion.
@@ 5. Path partial groups
As all these vertices belong to the clique $C$, this sequence can be thought of as the path ($(\{{v_{i_{1}}},v_{i_{2}}\},\ldots,\{v_{i_{r-1}},v_{i_{r}}\})$ in $G$. For $r=1$, we have the path consisting of a single vertex, $(v_{i_{1}})$. Thus elements in ${\mathcal{M}}$ can be seen as paths in the graph $G$ with $1$ or more vertices.
@@
where the two consecutive elements $b$ cancel each other as they correspond to the same element in ${\mathbb{Z}}_{2}$. Then $(a,c)$ should be a path in $G$ and the edge $\{a,c\}$ should be contained in $G$. Hence, paths that may be multiplied should have vertices contained in some clique of $G$.
@@
Since $\operatorname{Aut}({\mathbb{Z}}_{2})$ is trivial, the exact sequence (2) in Theorem 4.4 gives rise to an injective homomorphism $\Psi\colon\operatorname{Aut}_{{\mathcal{P}}art}({\mathbb{P}}(G))\to\operatorname{Aut}_{{\mathcal{G}}raphs}(G)$. In addition, the composition functor in Lemma 5.1 induces a group homomorphism
@@ 6. The graph of maximal locally finite subgroups of a partial group
Let ${\mathcal{X}}$ be the set of locally finite subgroups of $G$. Notice that ${\mathcal{X}}$ is not empty since $\{1\}$ is a locally finite subgroup of ${\mathbb{M}}$. If $H_{1}\leq H_{2}\leq\cdots$ is an increasing chain of locally finite subgroups of ${\mathbb{M}}$, consider the subgroup $H_{\infty}=\bigcup_{i\in{\mathbb{N}}}H_{i}\leq H$. By construction, every finitely generated subgroup of $H_{\infty}$ lies in some $H_{i}$ for $i$ big enough. Therefore, $H_{\infty}$ is a locally finite subgroup of $H$ and thus an upper bound for the given chain of elements of ${\mathcal{X}}$. Applying Zorn’s Lemma, we obtain that ${\mathcal{X}}$ has a maximal element. ∎
@@
Notice that the graph $\operatorname{MaxSub_{loc}}({\mathbb{M}})$ may be empty if ${\mathbb{M}}$ has no non-trivial subgroups, and Example 2.4 provides an example of such situation. In Definition 6.2, we look at locally finite subgroups instead of at subgroups because in a path partial group every edge give rise to a copy of ${\mathbb{Z}}$. This is demonstrated in Example 6.4 below.

### v6

@@ Abstract.
This work was partially supported by MCIN/AEI/10.13039/501100011033 [PID2020-118753GB-I00 to first and third authors, and PID2023-149804NB-I00 to third author].
@@ 1. Introduction
The proof of Theorem A is built upon a functor, defined in Section 3, that maps every pair $(G,\mathcal{H})$, for $G=(V,E)$ a simple undirected graph and =$\mathcal{H}=\{H_{v}\}_{v\in V}$ a collection of non-trivial groups, to a partial group ${\mathbb{M}}(G,\mathcal{H})$ such that the group of automorphisms $\operatorname{Aut}_{{\mathcal{P}}art}\big{(}{\mathbb{M}}(G,\mathcal{H})\big{)}$ is closely related to the groups $\operatorname{Aut}_{{\mathcal{G}}raphs}(G)$ and $\prod_{v\in V}\operatorname{Aut}_{{\mathcal{G}}roups}(H_{v})$ (see Theorem 4.4). The algebraic structure of the partial group ${\mathbb{M}}(G,\mathcal{H})$ is related to path concatenation within $G$. In fact, the special case with $\mathcal{H}$ consisting of copies of ${\mathbb{Z}}_{2}$, i.e., the cyclic group of order $2$, is the central character in this work: path partial groups.
@@
The algebraic structure of ${\mathbb{P}}(G)$ can be described in terms of the paths contained in the cliques of $G$ and path concatenation, together with the fact that vertices behave as inverses of themselves, see discussion after Lemma 5.1. Its construction resembles the ideas underlying path algebras, i.e., the algebra associated to a quiver [3, II.1.2]. The following result is the key ingredient in the proof of Theorem A.
@@ 2. Basics on partial groups
Let ${\mathcal{M}}$ be a non-empty set and $\mathbb{D}\subseteq\mathbb{W}({\mathcal{M}})$ be a subset such that,
@@
    if $u\in\mathbb{D}$ then $u^{-1}\circ u\in\mathbb{D}$ and $\Pi(u^{-1}\circ u)=1$.
A word $w\in\mathbb{D}$ will be called non-degenerate if there is no $1$ in $w$. Then, $\mathbb{D}$ is totally determined by its non-degenerate words.
@@
Let $\mathcal{F}(a)=\{1,a,a^{-1}\}$ and set the non-degenerate words of $\mathbb{D}_{a}$ to be all possible words alternating $a$’s and $a^{-1}$’s. In other words, the non-degenerate words of $\mathbb{D}_{a}$ are all the different finite sub-words of the infinite word $(a,a^{-1},a,a^{-1},a,a^{-1},\dots)$. The inversion is understood and, for any word $u\in\mathbb{D}\left(\mathcal{F}(a)\right)$,
@@
A homomorphism $\varphi\colon{\mathcal{M}}_{1}\to{\mathcal{M}}_{2}$ is an isomorphism of partial groups if the map $\varphi$ is bijective and $\varphi^{-1}$ is also a morphism of partial groups. Finally, an automorphism of the partial group ${\mathcal{M}}_{1}$ is an isomorphism $\varphi\colon{\mathcal{M}}_{1}\to{\mathcal{M}}_{1}$.
@@
has alphabet ${\mathcal{M}}$ equal to the quotient set
${\mathcal{M}}=\cup_{C\in\mathcal{C}}{\mathcal{M}}(C)/\sim.$
Here, $\sim$ is the smallest equivalence relation that contains the following relation,
@@
and such that there is a map $\Pi$ closing the following diagram,
@@
In this diagram, $\mathbb{D}$ is defined as the subset of words $u$ in the free monoid $\mathbb{W}({\mathcal{M}})$ for which there exists an object $C\in\operatorname{Ob}(\mathcal{C})$ and a word $v$ in $\mathbb{D}(C)$ such that $u$ is the image of $v$ under the component-wise application of $\sim$, the rightmost vertical arrow is induced by $\sim$, and the leftmost vertical arrow by its component-wise application.
@@ 3. Partial groups out of simple graphs
In what follows, $G=(V,E)$ is an object in ${{\mathcal{G}}raphs}$ and =$\mathcal{H}=\{H_{v}\}_{v\in V}$ is a collection of groups. Our aim is to define a partial group depending on $G$ and $\mathcal{H}$. Now, for a clique $C\subseteq V$ in $G$, we consider the free product $H_{C}=\mathop{\ast}\limits_{v\in C}H_{v}$ and the following construction.
@@
    $(u_{1},u_{2},\dots,u_{n})\in\mathbb{D}(C)\mapsto\Pi(u_{1}\circ u_{2}\circ\cdots\circ u_{n})\in\operatorname{CR}(C)\text{, and}$
@@
$(x_{1},x_{2},\dots,x_{2n}):=w\circ w^{-1}=(u_{1},u_{2},\dots,u_{n},u_{n}^{-1},u_{n-1}^{-1},\dots,u_{1}).$
@@
Let $\Delta$ be the poset of cliques $C\subseteq V$ in $G$. An inclusion $C_{1}\subseteq C_{2}$ between cliques gives rise to an inclusion map $\operatorname{CR}(C_{1})\subseteq\operatorname{CR}(C_{2})$ and to its associated map of partial groups ${\mathbb{M}}(C_{1})\to{\mathbb{M}}(C_{2})$. This way, we have a functor ${\mathbb{M}}(-)\colon\Delta\to{{\mathcal{P}}art}$.
@@
Then, from the simple shape of the category $\Delta$ and the functor ${\mathbb{M}}(-)$, together with uniqueness of the reduced form ($\Pi$), it follows that map
@@
is well defined and injective, where $(g_{1},g_{2},\ldots,g_{r})\in\operatorname{CR}(C)$ for some clique $C$. From now on, we will make the identification given by (S) and the corresponding one for $\mathbb{D}(G,\mathcal{H})$ without further notice.
@@ 4. Graph morphisms from partial groups homomorphisms
The category ${{\mathcal{I}}njPart}$ is the wide subcategory of the category ${{\mathcal{P}}art}$ with morphisms the partial group homomorphisms whose kernel has no torsion, i.e., it contains no non-trivial finite subgroup.
@@ 5. Path partial groups
As all these vertices belong to the clique $C$, this sequence can be thought of as the path ($(\{{v_{i_{1}}},v_{i_{2}}\},\ldots,\{v_{i_{r-1}},v_{i_{r}}\})$ in $G$. For $r=1$, we have the path consisting of a single vertex, $(v_{i_{1}})$. Thus elements in ${\mathcal{M}}$ can be seen as paths lying in some clique of $G$ and with $1$ or more vertices.
@@
where the two consecutive elements $b$ cancel each other as they correspond to the same element in ${\mathbb{Z}}_{2}$. Then $(a,c)$ should be a path in $G$ and the edge $\{a,c\}$ should be contained in $G$. Hence, paths that may be multiplied should have vertices contained in some clique of $G$ and, in addition, multiplication is path concatenation with the extra feature that two consecutive identical vertices produce the empty path.
@@
Since $\operatorname{Aut}({\mathbb{Z}}_{2})$ is trivial, the exact sequence (2) in Theorem 4.4 gives rise to an injective homomorphism $\Psi\colon\operatorname{Aut}_{{\mathcal{P}}art}({\mathbb{P}}(G))\to\operatorname{Aut}_{{\mathcal{G}}raphs}(G)$. In addition, the composite functor in Lemma 5.1 induces a group homomorphism
@@ 6. The graph of maximal locally finite subgroups of a partial group
Let ${\mathcal{X}}$ be the set of locally finite subgroups of ${\mathbb{M}}$. Notice that ${\mathcal{X}}$ is not empty since $\{1\}$ is a locally finite subgroup of ${\mathbb{M}}$. If $H_{1}\leq H_{2}\leq\cdots$ is an increasing chain of locally finite subgroups of ${\mathbb{M}}$, consider the subgroup $H_{\infty}=\bigcup_{i\in{\mathbb{N}}}H_{i}\leq{\mathcal{M}}$. By construction, every finitely generated subgroup of $H_{\infty}$ lies in some $H_{i}$ for $i$ big enough. Therefore, $H_{\infty}$ is a locally finite subgroup of ${\mathbb{M}}$ and thus an upper bound for the given chain of elements of ${\mathcal{X}}$. Applying Zorn’s Lemma, we obtain that ${\mathcal{X}}$ has a maximal element. ∎
@@
If ${\mathbb{M}}$ has no non-trivial subgroups, then the graph $\operatorname{MaxSub_{loc}}({\mathbb{M}})$ contains only the trivial subgroup, and Example 2.4 provides an example of such situation. In Definition 6.2, we look at locally finite subgroups instead of at subgroups because in a path partial group every edge give rise to a copy of ${\mathbb{Z}}$. This is demonstrated in Example 6.4 below.
@@
7. Declarations
The authors have no conflicts of interest to declare that are relevant to the content of this article.